In a survey of men in the United


States (ages 20-29), the mean height was 69.6 inches with a standard deviation of 3.0 inches. Find the minimum height in the top 16%.


1
Expert's answer
2021-06-03T07:01:25-0400

"\mu=69.6 \\

\sigma=3.0 \\

P(X>x) = 0.16 \\

P(X<x) = 1-0.16 = 0.84 \\

P(Z< \frac{x-\mu}{\sigma}) = 0.84 \\

\frac{x-69.6}{3.0}=0.995 \\

x-69.6= 3.0 \times 0.995 \\

x = 69.6+2.985 = 72.585"

Answer: 72.58 inches